# Generalization Bounds for Semi-supervised Matrix Completion with Distributional Side Information

### v1

@@ Abstract
We study a matrix completion problem where both the ground truth $R$ matrix and the unknown sampling distribution $P$ over observed entries are low-rank matrices, and share a common subspace. We assume that a large amount $M$ of unlabeled data drawn from the sampling distribution $P$ is available, together with a small amount $N$ of labeled data drawn from the same distribution and noisy estimates of the corresponding ground truth entries. This setting is inspired by recommender systems scenarios where the unlabeled data corresponds to ‘implicit feedback’ (consisting in interactions such as purchase, click, etc. ) and the labeled data corresponds to the ‘explicit feedback’, consisting of interactions where the user has given an explicit rating to the item. Leveraging powerful results from the theory of low-rank subspace recovery, together with classic generalization bounds for matrix completion models, we show error bounds consisting of a sum of two error terms scaling as $\widetilde{O}\left(\sqrt{\frac{nd}{M}}\right)$ and $\widetilde{O}\left(\sqrt{\frac{dr}{N}}\right)$ respectively, where $d$ is the rank of $P$ and $r$ is the rank of $M$. In synthetic experiments, we confirm that the true generalization error naturally splits into independent error terms corresponding to the estimations of $P$ and and the ground truth matrix $\operatorname{G}$ respectively. In real-life experiments on Douban and MovieLens with most explicit ratings removed, we demonstrate that the method can outperform baselines relying only on the explicit ratings, demonstrating that our assumptions provide a valid toy theoretical setting to study the interaction between explicit and implicit feedbacks in recommender systems.
@@ Introduction
  • We propose a semi-supervised learning paradigm for matrix completion: we assume that the sampling distribution $P\in[0,1]^{m\times n}$ over entries shares a low-rank subspace with the ground truth matrix $\operatorname{G}$, and that that a large number $M$ of unlabeled interactions is provided, together with a much smaller number $N$ of labeled interactions.
@@
where the notation $\widetilde{O}$ hides polylogarithmic factors in $m,n,\delta$. Although the condition (15) implies the number of labeled samples $N$ must pass a threshold which is linear in the size of the full matrix, this is a fixed threshold which doesn’t depend on the desired error level. In addition, this is only a worst-case scenario: one can alternatively assume that Assumption 7 holds with $\Gamma\simeq d$, also allowing the absorption of the higher order term $\sqrt{\frac{\Gamma[m+n]r}{MN}}$ into the others.
@@
To demonstrate the decomposition of the error into independent terms corresponding to the estimation errors of the labeled and unlabeled data respectively, we generated $\operatorname{G},P\in\mathbb{R}^{200\times 200}$ satisfying our assumptions with $d=r=4$. The detailed experimental setup can be found in the Appendix. For a broad range of values of both $N$ and $M$ ($M\in\{10000,20000,\ldots,100000\}$ and $N\in\{50,100,150,\ldots,1000\}$), we evaluate the average generalization error over $30$ independent runs. This range is selected because $N=100,M=100000$ results in perfect recovery up to a high decimal point. We compared two quantities:
@@
    where $1000$ an $100000$ are the maximum possible values for $N$ and $M$ respectively.
@@
Table 1: Performance in terms of Root Mean Squared Error (RMSE) across three datasets for varying values of $p$. Lower is better. The best results in each setting are presented in boldface.
@@
We observe that DAMC, which relies on the unsupervised information in the pure implicit feedback, significantly outperforms classic methods which rely only on the labeled observations in most situations. This indicates that a relationship between the sampling distribution and the ground truth matrix indeed exists, lending legitimacy to our theoretical learning setting. In particular, DAMC significantly outperform its counterpart Softimpute (which is the same algorithm without using side information) for all nonzero values of $p$. Most notably, we observe that for large values of $p$ such as $p=0.90$ and $p=0.95$, many of the baseline models relying only on explicit ratings are not able to perform much better than random, whilst the semi-supervised DAMC can still achieve consistently good performance.
@@ Conclusion
We introduced a new matrix completion learning setting where the sampling distribution and the ground truth matrix are both low-rank and share common row and column subspaces. This setting is inspired by the recommender systems application, where unlabeled interactions (‘implicit feedback’) are usually much more abundant than labeled interactions (‘explicit feedback’). Assuming access to a larger amount $M$ of unlabeled samples and a smaller number $N$ of labeled samples, we show generalization error bounds of the form $\widetilde{O}\left(\sqrt{\frac{[m+n]r}{M}}+\sqrt{\frac{dr}{N}}+\sqrt{\frac{\Gamma[m+n]r}{MN}}\right).$ When either $\Gamma$ (the ratio between the maximum and average sampling probability) is $O(d)$ or $N\geq\frac{m+n}{2}$, the higher-order term $\sqrt{\frac{\Gamma[m+n]r}{MN}}$ vanishes, demonstrating a disentanglement between two sources of error: the estimation of the shared low-dimensional subspaces relying on the unlabeled samples, and the estimation of the ground truth matrix. In particular, our results demonstrate the ground truth matrix can be recovered accurately even with a vanishingly small number of labeled interactions per row/column. On real data, we show that unlabeled samples can dramatically improve the performance of explicit feedback prediction methods, lending validity to our assumptions. In future work, it would be interesting to distill the real-data results into a true SVD and to tackle the difficulty of removing Assumption 7 or the uniform marginals assumption by imposing suitable modifications on the algorithm, or to attempt to derive optimistic bounds with a fast decay rate in $N$.
@@ II Identity matrix
$2\operatorname{\mathcal{B}}\frac{\log(2/\delta)}{3N}+25{\mathcal{P}^{*}}\ell\log\left(\frac{4[m+n]}{\delta}\right)\left[3\kappa_{1}\sqrt{\frac{[m+n]r}{M}}+\sqrt{\frac{[m+n]\Gamma r}{MN}}\right]$ (A.47)
@@ Appendix C Proof of the Main Results
$\operatorname{l}\left(X\underline{M}Y^{\top}\right)$ $\leq\widehat{\operatorname{l}}\left(X\underline{M}Y^{\top}\right)+2\frac{\operatorname{\mathcal{B}}\log(6/\delta)}{\sqrt{N}}+16\operatorname{\ell}{\mathcal{P}^{*}}^{2}\sqrt{\kappa_{1}\kappa_{2}}\log(2de)\sqrt{\frac{dr}{N}}$
@@
$\quad\quad\quad\quad\quad\quad+2\operatorname{\mathcal{B}}\frac{\log(6/\delta)}{3N}+25{\mathcal{P}^{*}}\ell\log\left(\frac{12[m+n]}{\delta}\right)\left[3\kappa_{1}\sqrt{\frac{[m+n]r}{M}}+\sqrt{\frac{[m+n]\Gamma r}{MN}}\right]$ (A.6)
$\leq\widehat{\operatorname{l}}\left(X\underline{M}Y^{\top}\right)+2\frac{\operatorname{\mathcal{B}}\log(6/\delta)}{\sqrt{N}}+8\operatorname{\ell}{\mathcal{P}^{*}}^{2}\sqrt{\kappa_{1}\kappa_{2}}\log(2de)\sqrt{\frac{dr}{N}}\left[1+\sqrt{\frac{1}{N}}\right]$
@@ Appendix E Further Experimental Details
First, the $m=n=200$ rows and columns of the matrix were each evenly partitioned into 4 random groups of 50 rows or columns. The true side information matrices $X^{*}$ and $Y^{*}$ were set to be the indicators of the row and column groups. In recommender systems applications, we may interpret the groups as genres (for movies) or customer types. Then, we generated a random rank ${\underline{M}^{*}}\in\mathbb{R}^{4\times 4}$ with i.i.d. $N(0,1)$ entries. The matrix was then normalized to ensure $\|{\underline{M}^{*}}\|_{\operatorname{Fr}}=4$, which corresponds to entries of size $O(1)$. The sampling distribution over entries $P$ was selected randomly as follows: each entry $P^{0}_{u,v}$ of a matrix $P^{0}\in\mathbb{R}^{10\times 10}$ was drawn form a uniform $U[0,1]$ distribution, and the initial matrix $P$ was set to $P=X^{*}P^{0}{Y^{*}}^{\top}$. Then, the the full matrix $P\in\mathbb{R}^{m\times m}=\mathbb{R}^{200\times 200}$ was normalized by dividing by $|P|_{1}$ to ensure a valid probability distribution. The ground truth matrix was set to $X^{*}{\underline{M}^{*}}{Y^{*}}\top$ and observed with some noise of standard deviation $0.05$. For each configuration of $M,N$, we performed 30 runs of the full pipeline of (1) generation of the sampling matrix $P$, (2) generation of the the $M$ unlabeled datapoints and the $N$ labeled data points, as well as (3) the training of the model (with the coefficient of nuclear norm regularization set to zero) and evaluation on a held out test set. We calculated the generalization error for each configuration for $M\in\{10000,20000,\ldots,100000\}$ and $N\in\{50,100,150,\ldots,1000\}$. The values were selected to ensure that nearly perfect performance for the largest values of $M$ and $N$.
@@ Appendix F Classic Results Required for the Proofs
Let $X_{1},\ldots,X_{N}$ be $N$ independent random variables such that $|X_{i}|\leq b$ almost surely. Let $v=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}(X_{i}^{2})$. With probability greater than $1-\delta$, we have
$\frac{1}{N}\sum_{i=1}^{N}X_{i}-\mathbb{E}(X_{i})\leq b\frac{\log\left(\frac{1}{\delta}\right)}{3N}+\sqrt{\frac{2v\log\left(\frac{1}{\delta}\right)}{N}}.$ (A.4)

### v2

@@ Abstract
We study a matrix completion problem where both the ground truth $R$ matrix and the unknown sampling distribution $P$ over observed entries are low-rank matrices, and share a common subspace. We assume that a large amount $M$ of unlabeled data drawn from the sampling distribution $P$ is available, together with a small amount NN of labeled data drawn from the same distribution and noisy estimates of the corresponding ground truth entries. This setting is inspired by recommender systems scenarios where the unlabeled data corresponds to ‘implicit feedback’ (consisting in interactions such as purchase, click, etc. ) and the labeled data corresponds to the ‘explicit feedback’, consisting of interactions where the user has given an explicit rating to the item. Leveraging powerful results from the theory of low-rank subspace recovery, together with classic generalization bounds for matrix completion models, we show error bounds consisting of a sum of two error terms scaling as $\widetilde{O}\left(\sqrt{\frac{nd}{M}}\right)$ and $\widetilde{O}\left(\sqrt{\frac{dr}{N}}\right)$ respectively, where $d$ is the rank of $P$ and $r$ is the rank of $M$. In synthetic experiments, we confirm that the true generalization error naturally splits into independent error terms corresponding to the estimations of $P$ and and the ground truth matrix $\operatorname{G}$ respectively. In real-life experiments on Douban and MovieLens with most explicit ratings removed, we demonstrate that the method can outperform baselines relying only on the explicit ratings, demonstrating that our assumptions provide a valid toy theoretical setting to study the interaction between explicit and implicit feedbacks in recommender systems.
@@ Introduction
  • We propose a semi-supervised learning paradigm for matrix completion: we assume that the sampling distribution $P\in[0,1]^{m\times n}$ over entries shares a low-rank subspace with the ground truth matrix $\operatorname{G}$, and that that a large number $M$ of unlabeled interactions is provided, together with a much smaller number NN of labeled interactions.
@@
where the notation $\widetilde{O}$ hides polylogarithmic factors in $m,n,\delta$. Although the condition (15) implies the number of labeled samples NN must pass a threshold which is linear in the size of the full matrix, this is a fixed threshold which doesn’t depend on the desired error level. In addition, this is only a worst-case scenario: one can alternatively assume that Assumption 7 holds with $\Gamma\simeq d$, also allowing the absorption of the higher order term $\sqrt{\frac{\Gamma[m+n]r}{MN}}$ into the others.
@@
To demonstrate the decomposition of the error into independent terms corresponding to the estimation errors of the labeled and unlabeled data respectively, we generated $\operatorname{G},P\in\mathbb{R}^{200\times 200}$ satisfying our assumptions with $d=r=4$. The detailed experimental setup can be found in the Appendix. For a broad range of values of both NN and $M$ ($M\in\{10000,20000,\ldots,100000\}$ and $N\in\{50,100,150,\ldots,1000\}$), we evaluate the average generalization error over $30$ independent runs. This range is selected because $N=100,M=100000$ results in perfect recovery up to a high decimal point. We compared two quantities:
@@
    where $1000$ an $100000$ are the maximum possible values for NN and $M$ respectively.
@@
Table 1: Performance in terms of Root Mean Squared Error (RMSE) across three datasets for varying values of pp. Lower is better. The best results in each setting are presented in boldface.
@@
We observe that DAMC, which relies on the unsupervised information in the pure implicit feedback, significantly outperforms classic methods which rely only on the labeled observations in most situations. This indicates that a relationship between the sampling distribution and the ground truth matrix indeed exists, lending legitimacy to our theoretical learning setting. In particular, DAMC significantly outperform its counterpart Softimpute (which is the same algorithm without using side information) for all nonzero values of pp. Most notably, we observe that for large values of $p$ such as $p=0.90$ and $p=0.95$, many of the baseline models relying only on explicit ratings are not able to perform much better than random, whilst the semi-supervised DAMC can still achieve consistently good performance.
@@ Conclusion
We introduced a new matrix completion learning setting where the sampling distribution and the ground truth matrix are both low-rank and share common row and column subspaces. This setting is inspired by the recommender systems application, where unlabeled interactions (‘implicit feedback’) are usually much more abundant than labeled interactions (‘explicit feedback’). Assuming access to a larger amount $M$ of unlabeled samples and a smaller number NN of labeled samples, we show generalization error bounds of the form $\widetilde{O}\left(\sqrt{\frac{[m+n]r}{M}}+\sqrt{\frac{dr}{N}}+\sqrt{\frac{\Gamma[m+n]r}{MN}}\right).$ When either $\Gamma$ (the ratio between the maximum and average sampling probability) is $O(d)$ or $N\geq\frac{m+n}{2}$, the higher-order term $\sqrt{\frac{\Gamma[m+n]r}{MN}}$ vanishes, demonstrating a disentanglement between two sources of error: the estimation of the shared low-dimensional subspaces relying on the unlabeled samples, and the estimation of the ground truth matrix. In particular, our results demonstrate the ground truth matrix can be recovered accurately even with a vanishingly small number of labeled interactions per row/column. On real data, we show that unlabeled samples can dramatically improve the performance of explicit feedback prediction methods, lending validity to our assumptions. In future work, it would be interesting to distill the real-data results into a true SVD and to tackle the difficulty of removing Assumption 7 or the uniform marginals assumption by imposing suitable modifications on the algorithm, or to attempt to derive optimistic bounds with a fast decay rate in NN.
@@ II Identity matrix
$4\operatorname{\mathcal{B}}\frac{\log(2/\delta)}{3N}+25{\mathcal{P}^{*}}\ell\log\left(\frac{4[m+n]}{\delta}\right)\left[3\kappa_{1}\sqrt{\frac{[m+n]r}{M}}+\sqrt{\frac{[m+n]\Gamma r}{MN}}\right]$ (A.47)
@@ Appendix C Proof of the Main Results
$\operatorname{l}\left(X\underline{M}Y^{\top}\right)$ $\leq\widehat{\operatorname{l}}\left(X\underline{M}Y^{\top}\right)+2.5\frac{\operatorname{\mathcal{B}}\log(6/\delta)}{\sqrt{N}}+16\operatorname{\ell}{\mathcal{P}^{*}}^{2}\sqrt{\kappa_{1}\kappa_{2}}\log(2de)\sqrt{\frac{dr}{N}}$
@@
$\quad\quad\quad\quad\quad\quad+4\operatorname{\mathcal{B}}\frac{\log(6/\delta)}{3N}+25{\mathcal{P}^{*}}\ell\log\left(\frac{12[m+n]}{\delta}\right)\left[3\kappa_{1}\sqrt{\frac{[m+n]r}{M}}+\sqrt{\frac{[m+n]\Gamma r}{MN}}\right]$ (A.6)
$\leq\widehat{\operatorname{l}}\left(X\underline{M}Y^{\top}\right)+2.5\frac{\operatorname{\mathcal{B}}\log(6/\delta)}{\sqrt{N}}+8\operatorname{\ell}{\mathcal{P}^{*}}^{2}\sqrt{\kappa_{1}\kappa_{2}}\log(2de)\sqrt{\frac{dr}{N}}\left[1+\sqrt{\frac{1}{N}}\right]$
@@ Appendix E Further Experimental Details
First, the $m=n=200$ rows and columns of the matrix were each evenly partitioned into 4 random groups of 50 rows or columns. The true side information matrices $X^{*}$ and $Y^{*}$ were set to be the indicators of the row and column groups. In recommender systems applications, we may interpret the groups as genres (for movies) or customer types. Then, we generated a random rank ${\underline{M}^{*}}\in\mathbb{R}^{4\times 4}$ with i.i.d. $N(0,1)$ entries. The matrix was then normalized to ensure $\|{\underline{M}^{*}}\|_{\operatorname{Fr}}=4$, which corresponds to entries of size $O(1)$. The sampling distribution over entries $P$ was selected randomly as follows: each entry $P^{0}_{u,v}$ of a matrix $P^{0}\in\mathbb{R}^{10\times 10}$ was drawn form a uniform $U[0,1]$ distribution, and the initial matrix $P$ was set to $P=X^{*}P^{0}{Y^{*}}^{\top}$. Then, the the full matrix $P\in\mathbb{R}^{m\times m}=\mathbb{R}^{200\times 200}$ was normalized by dividing by $|P|_{1}$ to ensure a valid probability distribution. The ground truth matrix was set to $X^{*}{\underline{M}^{*}}{Y^{*}}\top$ and observed with some noise of standard deviation $0.05$. For each configuration of $M,N$, we performed 30 runs of the full pipeline of (1) generation of the sampling matrix $P$, (2) generation of the the $M$ unlabeled datapoints and the NN labeled data points, as well as (3) the training of the model (with the coefficient of nuclear norm regularization set to zero) and evaluation on a held out test set. We calculated the generalization error for each configuration for $M\in\{10000,20000,\ldots,100000\}$ and $N\in\{50,100,150,\ldots,1000\}$. The values were selected to ensure that nearly perfect performance for the largest values of $M$ and NN.
@@ Appendix F Classic Results Required for the Proofs
Let $X_{1},\ldots,X_{N}$ be NN independent random variables such that $|X_{i}|\leq b$ almost surely. Let $v=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}(X_{i}^{2})$. With probability greater than $1-\delta$, we have
$\frac{1}{N}\sum_{i=1}^{N}X_{i}-\mathbb{E}(X_{i})\leq 2b\frac{\log\left(\frac{1}{\delta}\right)}{3N}+\sqrt{\frac{2v\log\left(\frac{1}{\delta}\right)}{N}}.$ (A.4)